Milo Clothing experienced the following events during 2018, its first year of operation:

  1. Acquired $18,500 cash from the issue of common stock.
  2. Purchased inventory for $6,500 cash.
  3. Sold inventory costing $3,900 for $6,630 cash.
  4. Paid $950 for advertising expense.

Required

Record the events in a horizontal statement model. In the Cash Flow column, use OA to designate operating activity, IA for investment activity, FA for financing activity, or NC for net change in cash. If the element is not affected by the event, leave the cell blank
Answer #1
MILO CLOTHING
Effect of Events on the Financial Statements
Event Assets = Stockholders' Equity Revenue - Expenses = Net Income Statement of Cash Flows
Cash + Inventory = Common Stock + Retained Earnings
1 $18500 + = $18500 + - = $18500 FA
2 -6500 + 6500 = + - = -6500 OA
3a 6630 + = + 6630 6630 - =

6630

6630 OA
3b + -3900 = + -3900 - 3900 = -3900
4 -950 + = + -950 - 950 = -950 -950 OA
Total $17680 + $2600 = $18500 + $1780 $6630 - $4850 = $1780 $17680 NC